OpenStudy (anonymous):

Step-by-step how do i factor y²+14s+49

OpenStudy (anonymous):

is it s in middle term

OpenStudy (anonymous):

THX mark o. :)

OpenStudy (zarkon):

@mark o. what happened to the y and s

myininaya (myininaya):

they became x x=y=s right mark-0 lol

OpenStudy (anonymous):

yeah same question i too have to ask

OpenStudy (anonymous):

i too posted (y+7)^2 but after seeing s i deleted it

OpenStudy (anonymous):

right myininaya ....thnx.....lol

OpenStudy (anonymous):

step by step:

OpenStudy (anonymous):

y\[y^2+14s+49 find the common factor\]

OpenStudy (anonymous):

wtf no such equation can be factorised

OpenStudy (anonymous):

but what common factor u will find fahminator

OpenStudy (anonymous):

it has to be 14y not 14s..... this cannot be factorised

OpenStudy (anonymous):

no common factor

OpenStudy (anonymous):

IM confused now :(

myininaya (myininaya):

do you realize you have s and y in your expression, joyce?

OpenStudy (anonymous):

therefore we had do multiply the first coefficient by the 3rd coefficient and find if i can use the factors of 49 to add to make 14

OpenStudy (anonymous):

yeah we know u r confused so after posting such question u thanked mark. o lol

OpenStudy (anonymous):

yeah you can’t do it if theres 14s... it has to be 14y OR s^2 in the beginning

myininaya (myininaya):

\[y^2+14y+49 \text{ can be factored}\] \[s^2+14s+49 \text{ can be factored}\] \[y^2+14s+49 \text{ cannot be factored}\]

OpenStudy (anonymous):

this equation was falsely written. this cannot be factorised any further

OpenStudy (anonymous):

tis is what i mean ..lol,,it should be y not s y^²+14y+49 =(y+7)(y+7) =(y+7)^2

OpenStudy (anonymous):

oh nvm then because its completely false using a new variable in the equation...

OpenStudy (anonymous):

i put down exactly wut the Q was

OpenStudy (anonymous):

lol then its impossible

myininaya (myininaya):

ok then it isn't factorable

OpenStudy (anonymous):

sometimes it cannot be factorable

OpenStudy (anonymous):

ANY FURTHER than what they have given you.

OpenStudy (anonymous):

ok thx

OpenStudy (anonymous):

just write down : impossible

OpenStudy (anonymous):

kk thx

OpenStudy (anonymous):

what about this problem x²+2x-15

myininaya (myininaya):

-15=5(-3) and 2=5+(-3)

myininaya (myininaya):

\[(x+5)(x-3)\]

OpenStudy (anonymous):

kk thx